Determine which of the following is not the solution set of the linear equations below.

3x-y+z=2

2x-z=2


1
Expert's answer
2022-02-14T16:29:01-0500

3x-y+z=2................(i)

2x-z=2...................(ii)

therefore z=2x-2

replace z in equation (i)

3x-y+2x-2=2

5x-y=4.......(iii)

the equation cannot be solved using substitution method or elimination since equation 2 lacks 3 variables